# Black Hole from Entropy Maximization

### v4

@@ VI.2.2 Exact saturation of a local sufficient condition by $g_{\mu\nu}^{*}$
10. Phenomenology—. We finally discuss the phenomenological aspect of the semi-classical gravity condensate. In Ref.CY , an investigation of imaging of the gravity condensate showed that, despite the absence of an event horizon, the image is significantly darkened by the strong redshift of (3) and almost identical to the classical black-hole image, giving the consistency with the current data. Furthermore, the intensity around the inner shadow is slightly enhanced when the emission is a bit inside the surface, which may be a future observable prediction for characterizing the condensate. It is also interesting to investigate gravitational waves in this model; in particular, some echo signal can be expected due to the existence of the surface structure echo . For a more realistic phenomenology, furthermore, it should be important to generalize the gravity condensate to a rotating case (see Ref.KY2 for a slowly-rotating case).
@@ Acknowledgments
Y.Y. thanks C.Barcelo, F.Becattini, R.Casadio, C.Y.Chen, C.Goeller, T.Harada, C.Kelly, E.Livine, N.Nakagawa, A.Pesci, and Y.Sakatani for inspiring discussions and valuable comments. Y.Y. is partially supported by Japan Society of Promotion of Science (Grants No.21K13929) and by RIKEN iTHEMS Program.
@@ A.2 Entropy
Now, let us evaluate the entropy. First, the equation of state $\rho=3p$ and thermodynamics lead to the Stefan-Boltzmann law:
@@ A.3 Consistency check
Third, we study the relation to the second law. Locally $\frac{\partial\tilde{s}(r)}{\partial T_{loc}(r)}>0$ holds from (74), while for the whole part, we have $\frac{\partial S}{\partial T_{0}}<0$ indicating that the total heat capacity is negative Landau_SM , where we use (75) and $T_{0}=T_{loc}(R)\propto R^{-1/2}$ (from (73)). Thus, due to the long-range nature of gravity, the thermodynamic behavior is different depending on the region being considered.
@@
We next examine (57). The right hand side becomes $\rho+p=\frac{4}{3}\rho=\frac{1}{14\pi Gr^{2}}$ from (69). For the left hand side, we have $\alpha(r)=\frac{\partial_{r}\log\sqrt{-g_{tt}(r)}}{\sqrt{g_{rr}(r)}}\sim\frac{1}{r}$ from (71) and obtain from (74) $\left|\frac{\hbar\alpha}{2\pi}\frac{s}{4\pi r^{2}}-\frac{\hbar}{2\pi}\frac{1}{\sqrt{g_{rr}}}\partial_{r}\left(\frac{s}{4\pi r^{2}}\right)\right|\sim\frac{n_{f}^{\frac{1}{4}}}{Gr^{2}}\sqrt{\frac{l_{p}}{r}}$. Thus, (57) is not saturated.
@@
We give a short review for the derivation of the self-consistent values (33) of $(\sigma,\eta)$ where the origin of the tangential pressure can be seen (see Ref.KY4 for details). This is also a demonstration of our self-consistent argument in the semi-classical Einstein equation (1).
We start with a review about how to solve the semi-classical Einstein equation (1) self-consistently. First, we consider the physical system and problem of interest and construct a candidate metric $g_{\mu\nu}$ for it, say, by making a model or a thought experiment. Next, using that metric as a background spacetime, we examine the behavior of the matter fields and identify a candidate state $|\psi\rangle$ for the system. Then, we use the solutions of the matter field equations, construct the regularized energy-momentum tensor, and renormalize it to remove divergences. Finally, we equate the obtained renormalized energy-momentum tensor $\langle\psi|T_{\mu\nu}|\psi\rangle$ with the Einstein tensor $G_{\mu\nu}$ calculated from the metric and solve (1). If it can be solved consistently, the candidate $(g_{\mu\nu},|\psi\rangle)$ is the self-consistent solution. If not, we set up another candidate and repeat the procedure. Note that this self-consistent analysis of (1) allows us to obtain a non-perturbative solution in $\hbar$.
Let’s implement this program by setting (32) for the inner region (28) and the Schwarzschild metric with $a_{0}$ for the outer one $R\leq r$ as the candidate metric.
@@
Here, $\alpha_{0}$ and $\beta_{0}$ fix a 4D theory at energy scale $\hbar\mu_{0}$ while we have chosen $\gamma_{0}=0$ because of the 4D Gauss-Bonnet theorem. Thus, we obtain a more explicit form of the renormalized energy-momentum tensor in the right hand side of (1):
@@
Using the fact that (32) is locally $AdS_{2}\times S^{2}$ (with AdS radius $L\equiv\sqrt{2\eta^{2}\sigma}$ and $S^{2}$ radius $r$), we can solve $\Box\phi=0$ around a point $r=r_{0}$ perturbatively by a $1/r$ expansion for $r\gg L$. Then, the 0-th order solution for the bound modes is given by
@@
Here, $\gamma$ is Euler’s constant and $c$ is the non-trivial finite value for $|0\rangle$: $c=0.055868$. We note that the trace part
@@
is independent of $(\alpha_{0},\beta_{0})$. Here, we can use (34) and check that this agrees with the leading value of the 4D Weyl anomaly BD ; Nicolai :
@@ Appendix C Maximum entropy from uniformity
We provide another derivation of (24): we derive it from a radially-uniform condition, instead of setting $\epsilon=\epsilon_{max}$. This indicates that the radial uniformity is a sufficient condition for entropy maximization; since the converse proposition of this is given in Sec.IV.1, we can therefore conclude that a necessary and sufficient condition for entropy maximization is radial uniformity. This would be natural for a spherically symmetric system in equilibrium according to thermodynamics in flat space, but it is non-trivial for a self-gravitating system.
@@
Comparing this to (17), we find $N_{0}\sim n$. Therefore, (107) means (24).
@@
We show that keeping Tolman’s law under the semi-classical condition (17) and consistency with local thermodynamics does not yield maximum entropy for a given surface area. This gives an explanation for the violation of Tolman’s law to obtain maximum entropy, discussed in Sec.V.1.1. Note that this is consistent with Refs.Green ; Xia , where isotropic fluid is assumed and quantities other than a surface area are fixed (see also footnote 9).
@@
where the maximum excitation $\epsilon_{max}\sim\frac{m_{p}}{\sqrt{n}}$ is reached at an innermost radius $r_{0}$. Then, (109) should saturate the local condition (18): $\frac{\hbar}{\epsilon(r)}\sim\sqrt{g_{rr}(r)}(r-a(r))=\frac{r}{\sqrt{g_{rr}(r)}}$, since it has led to (21). Taking the square of the both hands and applying (109), we get
@@
which is smaller than the maximum one (24) (and Buchdahl’s limit, $\frac{8}{9}r$ Buchdahl ). Through the right hand side of (21), therefore, the entropy of the configuration (115) is $\sim\frac{B_{0}-1}{B_{0}}\frac{R^{2}}{l_{p}^{2}}<\frac{2}{3}\frac{R^{2}}{l_{p}^{2}}$. This cannot exceed $\sim\frac{R^{2}}{l_{p}^{2}}$ (25), the value estimated for the gravity condensate in the same way. Thus, we conclude that Tolman’s law does not maximize entropy for a given surface area.
Finally, we discuss an interesting point emerging as a byproduct. For $B_{0}=2$, (116) gives $-G^{t}{}_{t}=G^{r}{}_{r}=G^{\theta}{}_{\theta}=\frac{1}{2r^{2}}$. This represents Zel’dovich’s causal-limit fluid, where $\rho=p$ holds for $\rho=\langle-T^{t}{}_{t}\rangle,~{}p=\langle T^{r}{}_{r}\rangle=\langle T^{\theta}{}_{\theta}\rangle$ Zeldovich . The entropy $S\propto\frac{R^{2}}{l_{p}^{2}}$ can be obtained directly by applying the Gibbs relation $T_{loc}s_{3d}=\rho+p$, Tolman’s law $T_{loc}(r)=\frac{T_{0}}{\sqrt{-g_{tt}(r)}}$ and the metric (115) to the formula (2). This is a result supporting the consistency of our typicality argument. Also, the area-scaling entropy is consistent with Ref.Banks . It would be interesting to study the relation between their derivation and ours.

### v5

@@ VI.2.2 Exact saturation of a local sufficient condition by $g_{\mu\nu}^{*}$
10. Phenomenology—. We finally discuss the phenomenological aspect of the gravity condensate. In Ref.CY , an investigation of imaging of the gravity condensate showed that, despite the absence of an event horizon, the image is significantly darkened by the strong redshift of (3) and almost identical to the classical black-hole image, giving the consistency with the current data. Furthermore, the intensity around the inner shadow is slightly enhanced when the emission is a bit inside the surface, which may be a future observable prediction for characterizing the condensate. It is also interesting to investigate gravitational waves in this model; in particular, some echo signal can be expected due to the existence of the surface structure echo . For a more realistic phenomenology, furthermore, it should be important to generalize the gravity condensate to a rotating case (see Ref.KY2 for a slowly-rotating case).
@@ Acknowledgments
Y.Y. thanks C.Barcelo, F.Becattini, R.Casadio, C.Y.Chen, C.Goeller, T.Harada, C.Kelly, E.Livine, N.Nakagawa, A.Pesci, and Y.Sakatani for inspiring discussions and valuable comments. Y.Y. is partially supported by Japan Society for the Promotion of Science (No.21K13929) and by RIKEN iTHEMS Program.
@@ A.2 Entropy
Now, let us evaluate the entropy. First, the equation of state $\rho=3p$ and (local) thermodynamics lead to the Stefan-Boltzmann law:
@@ A.3 Consistency check
Third, we study the relation to the second law. Locally $\frac{\partial\tilde{s}(r)}{\partial T_{loc}(r)}>0$ holds from (75), while for the whole part, we have $\frac{\partial S}{\partial T_{0}}<0$ indicating that the total specific heat is negative Landau_SM , where we use (76) and $T_{0}=T_{loc}(R)\propto R^{-1/2}$ (from (74)). Thus, due to the long-range nature of gravity, the thermodynamic behavior is different depending on the region being considered.
@@
We next examine (58). The right-hand side becomes $\rho+p=\frac{4}{3}\rho=\frac{1}{14\pi Gr^{2}}$ from (70). For the left hand side, we have $\alpha(r)=\frac{\partial_{r}\log\sqrt{-g_{tt}(r)}}{\sqrt{g_{rr}(r)}}\sim\frac{1}{r}$ from (72) and obtain from (75) $\left|\frac{\hbar\alpha}{2\pi}\frac{s}{4\pi r^{2}}-\frac{\hbar}{2\pi}\frac{1}{\sqrt{g_{rr}}}\partial_{r}\left(\frac{s}{4\pi r^{2}}\right)\right|\sim\frac{n_{f}^{\frac{1}{4}}}{Gr^{2}}\sqrt{\frac{l_{p}}{r}}$. Thus, (58) is not saturated.
@@
We give a short review for the derivation of the self-consistent values (33) of $(\sigma,\eta)$, where the origin of the tangential pressure can be seen (see Ref.KY4 for details). This is also a demonstration of our self-consistent treatment in the semi-classical Einstein equation (1).
We start with a review about how to solve the semi-classical Einstein equation (1) self-consistently. First, we consider the physical system and problem of interest and construct a candidate metric $g_{\mu\nu}$ for it, say, by making a model or a thought experiment. Next, using that metric as a background spacetime, we examine the behavior of the matter fields and identify a candidate state $|\psi\rangle$ for the system. Then, we use the solutions of the matter field equations, construct the regularized energy-momentum tensor, and renormalize it to remove the UV divergences. Finally, we equate the obtained renormalized energy-momentum tensor $\langle\psi|T_{\mu\nu}|\psi\rangle$ with the Einstein tensor $G_{\mu\nu}$ calculated from the metric and solve (1). If it can be solved consistently, the candidate $(g_{\mu\nu},|\psi\rangle)$ is the self-consistent solution. If not, we set up another candidate and repeat the procedure. Note that this self-consistent analysis of (1) allows us to obtain a non-perturbative solution in $\hbar$.
Let us implement this program by setting (32) for the inner region (28) and the Schwarzschild metric with $a_{0}$ for the outer one $R\leq r$ as the candidate metric.
@@
Here, $\alpha_{0}$ and $\beta_{0}$ fix a 4D theory at energy scale $\hbar\mu_{0}$ while we have chosen $\gamma_{0}=0$ because of the 4D Gauss-Bonnet theorem. Thus, we obtain a more explicit form of the renormalized energy-momentum tensor in the right-hand side of (1):
@@
Using the fact that (32) is locally $AdS_{2}\times S^{2}$ (with AdS radius $L\equiv\sqrt{2\eta^{2}\sigma}$ and $S^{2}$ radius $r$ as in footnote 11), we can solve $\Box\phi=0$ around a point $r=r_{0}$ perturbatively by a $1/r$ expansion for $r\gg L$. Then, the 0-th order solution for the bound modes is given by
@@
Here, $\gamma$ is Euler’s constant and $c$ is a non-trivial finite value for $|0\rangle$: $c=0.055868$. We note that the trace part
@@
is independent of $(\alpha_{0},\beta_{0})$. We can use (34) and check that this agrees with the leading value of the 4D Weyl anomaly BD ; Nicolai :
@@ Appendix C Maximum entropy from uniformity
We provide another derivation of (24): we derive it from a radially-uniform condition, instead of setting $\epsilon=\epsilon_{max}$. This indicates that the radial uniformity is a sufficient condition for entropy maximization; since the converse proposition of this is given in Sec.IV.1, we can therefore conclude that a necessary and sufficient condition for entropy maximization is radial uniformity. This would be natural for a spherically symmetric system in equilibrium according to thermodynamics in flat space, but it is non-trivial for a self-gravitating system and consistent with Appendix.D.
@@
Comparing this to (17), we can identify $N_{0}\sim n$, since at this stage, $n$ is an arbitrary large number with $\mathcal{O}(1)$. Therefore, (108) means (24).
@@
We show that keeping Tolman’s law under the semi-classical condition (17) and consistency with local thermodynamics does not yield maximum entropy for a given surface area. This gives an explanation for the violation of Tolman’s law in maximizing the entropy (discussed in Sec.V.1.1). Note that this is consistent with Refs.Green ; Xia , where isotropic fluid is assumed and quantities other than a surface area are fixed (see also footnote 9).
@@
where the maximum excitation $\epsilon_{max}\sim\frac{m_{p}}{\sqrt{n}}$ is reached at an innermost radius $r_{0}$. To maximize the entropy in the bound (21), then (110) should saturate the local condition (18): $\frac{\hbar}{\epsilon(r)}\sim\sqrt{g_{rr}(r)}(r-a(r))=\frac{r}{\sqrt{g_{rr}(r)}}$. Taking the square of the both hands and applying (110), we get
@@
which is smaller than the maximum one (24) (and Buchdahl’s limit, $\frac{8}{9}r$ Buchdahl ). Through the right-hand side of (21), therefore, the entropy of the configuration (116) is $\sim\frac{B_{0}-1}{B_{0}}\frac{R^{2}}{l_{p}^{2}}<\frac{2}{3}\frac{R^{2}}{l_{p}^{2}}$. This cannot exceed $\sim\frac{R^{2}}{l_{p}^{2}}$ (25), the value estimated for the gravity condensate in the same way. Thus, we conclude that Tolman’s law does not maximize entropy for a given surface area.
Finally, we discuss an interesting point emerging as a by-product. For $B_{0}=2$, (117) gives $-G^{t}{}_{t}=G^{r}{}_{r}=G^{\theta}{}_{\theta}=\frac{1}{2r^{2}}$. This represents Zel’dovich’s causal-limit fluid, where $\rho=p$ holds for $\rho=\langle-T^{t}{}_{t}\rangle,~{}p=\langle T^{r}{}_{r}\rangle=\langle T^{\theta}{}_{\theta}\rangle$ Zeldovich . The entropy $S\propto\frac{R^{2}}{l_{p}^{2}}$ can be obtained directly by applying the Gibbs relation $T_{loc}s_{3d}=\rho+p$, Tolman’s law $T_{loc}(r)=\frac{T_{0}}{\sqrt{-g_{tt}(r)}}$ and the metric (116) to the formula (2). This is a result supporting the consistency of our typicality argument. Also, the area-scaling entropy is consistent with Ref.Banks . It would be interesting to study the relation between their derivation and ours.